Controls Meeting 7-Mar-2013
- Permanent room change to A110
- Josh gone, Dave in charge, Scot training to be backup
- Minutes from previous meeting 7-Feb-2013
- FCAL Controls Update - Dan, John
- Update on PLC test stand - Dave
- Counting House computer system - Hovanes, Elliott
- UPS power in the hall - Elliott
- EPICS archiver update - Hovanes
- Solenoid test status - Josh, Mark, Scot, Elliott
- SOE, PXI and UPS, Danfysik fan and UPS, QD, EWEB
- Other controls applications - Dave
- JInventory - Elliott
- Mantis task tracker
- Midas elog
- Experiment Control Supervisor - Elliott
Note permanent room change!
9:30AM Thurs 7-Mar-2013 CC A110
9:30am Thurs 21-Mar-2013
Present: Elliott W, Dave B, Scot S, Dan B, John L.
FCAL Controls Update
Dan and John presented their latest thoughts on the FCAL darkroom structure and control system design (see slides above). New concerns were fire suppression and including a a kick-out panel. Still being worked out is choice of sensors for door positons, crash buttons, required signage, who will do the wiring and where, documentation of the control system wiring and a number of other items. They will produce a list of required sensors and Scot and Mark will make recommendations based on JLab standard practice.
Tentative decision was to have IU personnel do the wiring here with guidance on JLab standard practices from Scot and Mark, and that JLab personnel will install the PLC and connect the IU wiring to it.
Controls Test Stand
Dave reported that the test stand is working well but that it needs to be expanded. He plans to order another desktop rack to hold more equipment.
Counting House Computers
Elliott reported that the initial set of computers, servers, switches, etc. is installed and mostly working. In particular the file server is up, accounts are being created as we speak, and disk partitions and areas are being configured. We plan to purchase additional equipment to replace the borrowed stuff in conjunction with the large equipment order by the Computer div. planned for late Spring.
UPS Power in the Hall
Elliott reported that current (tentative) plans are to run filtered power from the Rack room UPS to the hall and to install branch circuits to locations where UPS power is needed. The alternative, purchasing many small FCC Class B UPS's and maintaining them individually, is thought to be less cost effective. This work will not happen for quite some time so we will use small FCC Class A (existing) UPS's in the interim.
Note that we have no plans to run UPS power from the Rack room to the Tagger building, so we'll have to purchase Class B UPS's for equipment there.
Elliott reported that we have come to agreement with the Accelerator div. to purchase about $21k worth of computer hardware so that they can implement their archiver (MyA) according to Hall D requirements. The computers would be located in their area and managed by them for us. Since our archiving requirements will be modest for quite a while it is not clear when the computers would be purchased.
Elliott reported that we are still in a holding pattern waiting for the Cryo system to come online. Tuning of the quench detector needs to be performed, as does recommissioning of the Danfysik power supply. Some problems with the control board are being investigated (later...board was fine, problem was with power transistors).
Other Controls Applications
Dave reported our understanding of requirements for other control systems in the hall range from excellent to being almost clueless. The sooner the detector groups get us requirements, the better. Elliott will try to pry this information from them.